OpenStudy (18jonea):

Solve |x + 9| = 8. @pooja195

pooja195 (pooja195):

Absolute value can either have a negative solution or a positive solution x+9=18 x+9=-18 Solve both for x you should get 2 answers :)

imqwerty (imqwerty):

:O

OpenStudy (18jonea):

18?

OpenStudy (18jonea):

or 8?

imqwerty (imqwerty):

how did u get that :)

OpenStudy (18jonea):

x=-1 or x=17

OpenStudy (anonymous):

X= 17,1

OpenStudy (18jonea):

but i was wondering why she put 18

imqwerty (imqwerty):

yea

imqwerty (imqwerty):

:) a typo

pooja195 (pooja195):

Sorry thats my fault >.< x+9=8-9=-1 x=-1 x+9=-8=9=-17 x=-17 So yes you are correct sorry about the mistake!